OpenStudy (bananas1234):

Arrange the following fractions in order from least to greatest. 3/5, 5/8, 6/7

OpenStudy (butterflydreamer):

you could approach this question by changing each fraction to a percentage :) Then compare them and arrange them by using their percentage

OpenStudy (rational):

( 3/5, 5/8, 6/7) = (0.6, 0.63, 0.86) Now compare

OpenStudy (bananas1234):

6/7,5/8,3/5?

OpenStudy (butterflydreamer):

the question tells you to arrange it from SMALLEST to largest. I think you went the wrong way around :P

OpenStudy (bananas1234):

oh oops i thought it was greatest to least because thats what my last question was.

OpenStudy (butterflydreamer):

oh haha. No problem ^_^ Just fix it up and you're all set!

OpenStudy (bananas1234):

thanks, so 3/5, 5/8, 6/7

OpenStudy (butterflydreamer):

yeppp!

OpenStudy (bananas1234):

Thanks!

OpenStudy (butterflydreamer):

no worrriesssss :D

OpenStudy (dan815):

you should put them all over the same denominator if u ever wanna comapre fractions

OpenStudy (rational):

putting them in same denominator kinda takes same effort as long division

OpenStudy (bananas1234):

oh, so find what they have in common?

OpenStudy (dan815):

for example ( 3/5, 5/8, 6/7) = (0.6, 0.63, 0.86) = 60/100,63/100,86/100

OpenStudy (dan815):

anything over the same denominator we can compare we can compare 0.6,0.63,0.86 because they all have 1 as denominator u just dont write it out

OpenStudy (bananas1234):

is there more than one way to approach this?

OpenStudy (rational):

usually there will many ways to approach problems like these

OpenStudy (bananas1234):

so is turning them into decimals then comparing ok?

OpenStudy (rational):

If you're not a fan of long division, you may try multiplying all the fractions by 5*8*7 and compare the integers

OpenStudy (rational):

\[\dfrac{3}{5}~~\color{red}{?}~~ \dfrac{5}{8}~~\color{red}{?}~~\dfrac{6}{7}\] \[\color{blue}{(5*8*7)}\dfrac{3}{5}~~\color{red}{?}~~ \color{blue}{(5*8*7)}\dfrac{5}{8}~~\color{red}{?}~~\color{blue}{(5*8*7)}\dfrac{6}{7}\] \[\color{blue}{(8*7)}*3~~\color{red}{?}~~ \color{blue}{(5*7)}*5~~\color{red}{?}~~\color{blue}{(5*8)}*6\] \[168~~\color{red}{<}~~ 175~~\color{red}{<}~~240\]

OpenStudy (bananas1234):

i see that, that is a good way.